# First detections of PN, PO and PO+ toward a shocked low-mass starless core

### v1

@@ Abstract
Phosphorus is a key element that plays an essential role in biological processes important for living organisms on Earth. The origin and connection of phosphorus-bearing molecules to early Solar system objects and star-forming molecular clouds is therefore of great interest, yet there are limited observations throughout different stages of low-mass ($M<$ a few M) star formation. Observations from the Yebes 40 m and IRAM 30 m telescopes detect for the first time in the 7mm, 3mm, and 2mm bands multiple transitions of PN and PO, as well as a single transition of PO+, toward a low-mass starless core. The presence of PN, PO and PO+ is kinematically correlated with bright SiO(1-0) emission. Our results reveal not only that shocks are the main driver of releasing phosphorus from dust grains and into the gas-phase, but that the emission originates from gas not affiliated with the shock itself, but quiescent gas that has been shocked in the recent past. From radiative transfer calculations, the PO/PN abundance ratio is found to be $3.1^{+0.4}_{-0.6}$, consistent with other high-mass and low-mass star-forming regions. This first detection of PO+ toward any low-mass star-forming region reveals a PO+/PO ratio of $0.0115^{+0.0008}_{-0.0009}$, a factor of ten lower than previously determined from observations of a Galactic Center molecular cloud, suggesting its formation can occur under more standard Galactic cosmic-ray ionization rates. These results motivate the need for additional observations that can better disentangle the physical mechanisms and chemical drivers of this precursor of prebiotic chemistry.
@@ 1 Introduction
Although PN has been known for decades to be present in shells of evolved stars (Cernicharo et al., 2000; Tenenbaum et al., 2007) as well as massive star-forming regions (Turner & Bally, 1987; Ziurys, 1987; Turner et al., 1990), there are limited constraints for PN (as well as PO and PO+) in low-mass ($M<$ a few M) star-forming environments, with only a handful of detections toward low-mass protostars and outflows (Yamaguchi et al., 2011; Lefloch et al., 2016; Bergner et al., 2019; Wurmser & Bergner, 2022). We do know, however, that PN is present in quiescent high-mass dense cores (Fontani et al., 2016), as well as the central star-forming region of the distant starburst galaxy NGC 253 (Haasler et al., 2022). Still, there are no published constraints for PN, PO or the PO+ ion in the earlier, nearby ($<200$ pc), cold ($\sim$ 10 K), and moderately dense ($n\sim 10^{5}$ cm-3) low-mass starless and prestellar core stage.
@@
At a distance of 294 pc (see Pezzuto et al. 2021), the low-mass star-forming NGC 1333 region within the Perseus Molecular Cloud is known to host quiescent starless and prestellar cores as well as active protostars, which are associated with dynamic CO outflows and extended SiO emission red(Lefloch et al., 1998; Codella et al., 1999; Knee & Sandell, 2000; Plunkett et al., 2013). The widespread ($\sim 0.4$ pc across) and abundant SiO emission mapped by Lefloch et al. (1998) shows a narrow line ($\Delta v=0.4-1.2$ km/s) ‘ambient component’, separate from additional broad-line features associated with the high-velocity protostellar outflows, believed to be tracing rather the interaction of these outflows in the region with the dense gas. Codella et al. (1999) instead attributed the narrow-line SiO to fossil shocks from the protstellar outflows, or slower shocks that would require a significant amount of silicon on dust grain mantles. More recently, De Simone et al. (2022) hypothesized that this widespread narrow SiO is from expanding bubbles, or ‘large-scale turbulent cells’ (Dhabal et al., 2019), that collided with the cloud and triggered successive shocks. The the origin of these bubbles, however, is less clear.
Directly coincident with this ambient SiO emission in NGC 1333 is a ‘prestellar’ core $\sim$6200 au or ${}^{{}^{\prime\prime}}$ in size, labeled as such and indexed ‘326’ in the Herschel Perseus core catalog of Pezzuto et al. (2021) based firstly on the fact that no internal source of energy (e.g., protostar) is present, and secondly, on a virial analysis that found the core’s self-gravity exceeds its pressure support. We stress, however, that because the dynamical properties and other evolutionary parameters (e.g., infall signatures and a CO depletion factor) are not known, we will continue to refer to ‘Pers326’ as starless, not prestellar. The coordinates of Pers326 also coincides with an ammonia (NH3) peak (Rosolowsky et al., 2008; Kerr et al., 2019) and thus provides additional information such as the core’s kinetic temperature ($T_{\mathrm{k}}\sim 12$ K).
This criterion outlined above (Herschel plus NH3 observations) led to the investigation of the complex chemical inventory of Pers326. It was included in a sample of 35 total starless and prestellar cores in the Perseus Molecular Cloud surveyed with the Arizona Radio Observatory (ARO) 12 m dish, presenting high abundances of the simple COMs CH3OH and CH3CHO (Scibelli et al., 2024). Follow-up observations were done with the Yebes Observatory 40 m telescope’s Q-band receiver to search for higher complexity COMs in a sub-sample of the 15 cores with bright CH3OH and CH3CHO, including Pers326 (also presented in Scibelli et al. 2024). The wide-band nature of the Q-band observations led to the serendipitous detection of the PN(1-0) line, concurrent with bright SiO(1-0) emission, which we present here and confirm with additional observations at higher frequencies.
@@ 2 Observations
For both the Yebes 40 m and IRAM 30 m datasets the reduction was done with the publicly available Python-based scripts [^1] https://github.com/andresmegias/gildas-class-python/ developed by Megías et al. (2023), which invokes the CLASS program of the GILDAS package (Pety 2005; Gildas Team 2013) for the combining of the spectra (again, see Scibelli et al. 2024 for more detail). Data were inspected, baselined combined, and put on the main beam temperature, $T_{\mathrm{mb}}$, scale based on the public efficiency measurements given for each respective telescope and receiver. Gaussian fitting of spectral lines was done with the Python ‘Pyspeckit’ package (Ginsburg & Mirocha, 2011; Ginsburg et al., 2022). The lines presented and analyzed here are listed in Table 1 and plotted in Figures 2, 3, 4 and 5.
@@ 3 Detections and Line Properties
The wide-bandwidth in the Yebes 40 m data led to the serendipitous detection of the PN(1-0) line at 46.990 GHz toward the Pers326 starless core. A total of 15 cores bright in CH3OH and CH3CHO were observed in this sample (Scibelli et al., 2024), and so PN was also searched for in these cores with no success (see Table 2). Pers326 is the only core with strong ($T_{\mathrm{mb}}$ $>$ 1 K) SiO(1-0) emission at 43.423855 GHz [^2] Note that if one were to use the measured frequency of 43.42376 GHz from the JPL line list within the SPLATALOGUE database, SiO would be blue-shifted from the PN and PO emission by $\sim$0.8 km/s. This JPL entry has a roughly order of magnitude larger error than the entry in SLAIM, i.e., 50 kHz versus 3 kHz, respectively. One should thus be careful when extracting rest frequencies from SPLATALOGUE, and use those entries with smaller associated errors.  (Figure 2). This emission is only slightly blue-shifted, to 7.6 km/s, from the PN(1-0) $v_{\mathrm{lsr}}$ of 7.8 km/s, but given the coarse resolution of Yebes at these frequencies ($\sim$0.2 km/s) we conclude they are consistent (Table 1). A few other starless and prestellar cores reside nearby spatially to Pers326 with SiO detections, Pers264, Pers317 and Pers321 (see Figure 1), but the emission is much weaker ($T_{\mathrm{mb}}$ $<$ 0.3 K) and offset in $v_{\mathrm{lsr}}$ (Figure 2). For Pers264 the SiO(1-0) is blue-shifted more dramatically, to $\sim$2 km/s, suggesting this emission is not associated with the core itself (at a similar $v_{\mathrm{lsr}}$ to Pers326) but larger-scale surrounding SiO gas (see Figure 2). Given the strong SiO(1-0) and PN(1-0) correlation in Pers326, we suggest the origin of this gas-phase PN is related to shock chemistry (see section 5 for more discussion).
@@
Figure 2: SiO(1-0) emission observed toward Pers326 with the Yebes 40 m telescope (black) and a Gaussian fit to the line (blue). The vertical dashed lines correspond to the $v_{\mathrm{lsr}}$ of 7.6 km/s. Additional SiO emission toward the other nearby starless cores Pers321 (orange), Pers317 (red) and Pers264 (grey) are also plotted for reference.
@@
Figure 3: PN emission centered on the $F$ = 2-1 hyperfine component for each of the $J$ transitions observed toward Pers326 with both the Yebes 40 m and IRAM 30 m. As solid and blue and red curves are the Gaussian fits to the lines, the vertical dashed line shows a $v_{\mathrm{lsr}}$ of 7.6 km/s, and the horizontal grey line shows $1\sigma$ noise level. Spectra are offset by 100 mK for easier viewing. Annotated on the spectra are also the locations of the weaker and blended hyperfine components.
@@
Figure 4: PO emission observed toward Pers326 with the IRAM 30 m using both the 3 mm and 2 mm receivers. In solid red are the Gaussian fits to the detected ($>3\sigma$) lines and in dotted red are LTE fits at $T_{\mathrm{ex}}$ = 4.6 K. The vertical dashed line shows the consistent $v_{\mathrm{lsr}}$ of 7.6 km/s, and the horizontal grey line shows $1\sigma$ noise level. Spectra are offset by 80 mK for easier viewing. Full quantum numbers for each transition are given in Table 1.
@@
Figure 5: PO+ transitions observed toward Pers326. In solid red is the Gaussian fit to the detected ($>3\sigma$) line and the dotted lines are LTE fits at $T_{\mathrm{ex}}$ = 4.6 K. The vertical dashed line shows the consistent $v_{\mathrm{lsr}}$ of 7.6 km/s, and the horizontal grey line shows $1\sigma$ noise level. Spectra are offset by 40 mK for easier viewing.
@@ 4 Column Densities
Figure 6 shows, firstly, a least-squares fit to the PN data where $T_{\mathrm{ex}}=7.6\pm$0.7 K and $N_{\mathrm{tot}}=0.84\pm 0.12\times 10^{12}$ cm-2. A separate fit to the PO data finds a steeper slope where $T_{\mathrm{ex}}=4.63\pm$0.13 K and $N_{\mathrm{tot}}=2.63\pm 0.16\times 10^{12}$ cm-2. Uncertainties in the fits were calculated based on MonteCarlo simulations sampling the data points as Gaussians, and then analyzing the resulting distributions of the parameters using the richvalues Python package [^3] https://github.com/andresmegias/richvalues. While the spread in $E_{\mathrm{u}}$ for PO is limited, with only two distinct values (8.4 K and 15.7 K), we also fit synthetic spectra to each individual hyperfine line to find the parameters derived from the rotation diagram fit agree well with the data (see Figure 4). Similarly, using the calculated $T_{\mathrm{ex}}$ value of 4.6 K for PO, we fit synthetic spectra to the PO+ transitions and find a column density roughly two orders of magnitude lower than for PO, at $N_{\mathrm{tot}}=3.02^{+0.15}_{-0.12}\times 10^{10}$ cm-2. All column densities are reported in Table 3.
As seen before in the literature (e.g., Fontani et al. 2016; Wurmser & Bergner 2022) this emission is expected to be sub-thermally excited. Future work to model these P-bearing species in the optically thick limit and with non-LTE codes in conjunction with a better understanding of the physical structure (e.g., volume density) of the emitting gas can shed light on this, but is out of the scope for this current paper.
Additionally, we note that it is still unclear if the PN, PO and PO+ emission comes from a compact region around the core or if they trace the larger-scale SiO emission. Therefore, we assume the emission fills our beams ($f=1$) in all cases and thus our $N$ values are likely lower limits. We do note that for PN if we assume a smaller filling fraction (using the same source size of 51.5 arcsec as found for methanol in Scibelli et al. 2024), the $T_{\mathrm{ex}}$ drops to 6.1 K and the total column density is slightly raised, at $N_{\mathrm{tot}}=1.2\times 10^{12}$ cm-2.
@@ 5 Results and Discussion
We report the first detection of phosphorus-bearing molecules toward a shocked low-mass starless core. These observations provide new insight into the interstellar phosphorus budget in the earliest stage of star formation.
@@ 6 Conclusions
We thank the anonymous reviewer for their informative comments and feedback. The authors also thank Anthony Remijan and Víctor Rivilla for helpful discussion and suggestions. S.S. acknowledges the National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. I.J-.S and A.M. acknowledge funding from grant PID2022-136814NB-I00 funded by the Spanish Ministry of Science, Innovation and Universities/State Agency of Research MICIU/AEI/ 10.13039/501100011033 and by “ERDF/EU”, and from the ERC grant OPENS (project number 101125858) funded by the European Union. Yancy Shirley was supported by National Science foundation Astronomy and Astrophysics Grant AST-2205474. R.T.G. thanks the National Science Foundation for funding through the Astronomy & Astrophysics program (grant number 2206516). Observations carried out with the Yebes 40 m telescope (22A022 and 23A025). Paula Tarrío and Alba Vidal García carried out the observations and the first inspection of the data quality. The 40 m radio telescope at Yebes Observatory is operated by the Spanish Geographic Institute (IGN; Ministerio de Transportes, Movilidad y Agenda Urbana). Yebes Observatory thanks the European Union’s Horizon 2020 research and innovation programme for funding support to ORP project under grant agreement No 101004719. This work is also based on observations carried out under project number 024-24 with the IRAM 30 m telescope. IRAM is supported by INSU/CNRS (France), MPG (Germany) and IGN (Spain).

### v2

# First Detections of PN, PO, and PO+ toward a Shocked Low-mass Starless Core

@@ Abstract
Phosphorus is a key element that plays an essential role in biological processes important for living organisms on Earth. The origin and connection of phosphorus-bearing molecules to early Solar system objects and star-forming molecular clouds is therefore of great interest, yet there are limited observations throughout different stages of low-mass ($M<$ a few M) star formation. Observations from the Yebes 40 m and IRAM 30 m telescopes detect for the first time in the 7mm, 3mm, and 2mm bands multiple transitions of PN and PO, as well as a single transition of PO+, toward a low-mass starless core. The presence of PN, PO and PO+ is kinematically correlated with bright SiO(1-0) emission. Our results reveal not only that shocks are the main driver of releasing phosphorus from dust grains and into the gas-phase, but that the emission originates from gas not affiliated with the shock itself, but quiescent gas that has been shocked in the recent past. From radiative transfer calculations, the PO/PN abundance ratio is found to be $3.1^{+0.4}_{-0.6}$, consistent with other high-mass and low-mass star-forming regions. This first detection of PO+ toward any low-mass star-forming region reveals a PO+/PO ratio of $0.0115^{+0.0008}_{-0.0009}$, a factor of ten lowerk than previously determined from observations of a Galactic Center molecular cloud, suggesting its formation can occur under more standard Galactic cosmic-ray ionization rates. These results motivate the need for additional observations that can better disentangle the physical mechanisms and chemical drivers of this precursor of prebiotic chemistry.
@@ 1 Introduction
Although PN has been known for decades to be present in shells of evolved stars (Cernicharo et al., 2000; Tenenbaum et al., 2007) as well as massive star-forming regions (Turner & Bally, 1987; Ziurys, 1987; Turner et al., 1990), there are limited constraints for PN (as well as PO and PO+) in low-mass ($M<$ a few M) star-forming environments, with only a handful of detections toward low-mass protostars and outflows (Yamaguchi et al., 2011; Lefloch et al., 2016; Bergner et al., 2019; Wurmser & Bergner, 2022). We do know, however, that PN is present in quiescent high-mass dense cores (Fontani et al., 2016), as well as the central star-forming region of the distant starburst galaxy NGC 253 (Haasler et al., 2022). Still, there are no published constraints for PN, PO or the PO+ ion in the earlier, nearby ($<400$ pc), cold ($\sim$ 10 K), and moderately dense ($n\sim 10^{5}$ cm-3) low-mass starless and prestellar core stages.
@@
At a distance of 294 pc (see Pezzuto et al. 2021), the low-mass star-forming NGC 1333 region within the Perseus Molecular Cloud is known to host quiescent starless and prestellar cores as well as active protostars, which are associated with dynamic CO outflows and extended SiO emission red(Lefloch et al., 1998; Codella et al., 1999; Knee & Sandell, 2000; Plunkett et al., 2013). The widespread ($\sim 0.4$ pc across) and abundant SiO emission mapped by Lefloch et al. (1998) shows a narrow line ($\Delta v=0.4-1.2$ km/s) ‘ambient component’, separate from additional broad-line features associated with the high-velocity protostellar outflows, believed to be tracing rather the interaction of these outflows in the region with the dense gas. Codella et al. (1999) instead attributed the narrow-line SiO to fossil shocks from the protostellar outflows, or slower shocks that would require a significant amount of silicon on dust grain mantles. More recently, De Simone et al. (2022) hypothesized that this widespread narrow SiO is from expanding bubbles, or ‘large-scale turbulent cells’ (Dhabal et al., 2019), that collided with the cloud and triggered successive shocks. The origin of these bubbles, however, is less clear.
Directly coincident with this ambient SiO emission in NGC 1333 is a ‘prestellar’ core $\sim$6200 au or ${}^{{}^{\prime\prime}}$ in size, labeled as such and indexed ‘326’ in the Herschel Perseus core catalog of Pezzuto et al. (2021) based firstly on the fact that no internal source of energy (e.g., protostar) is present and second, on a virial analysis that found the core’s self-gravity exceeds its pressure support. We stress, however, that because the dynamical properties and other evolutionary parameters (e.g., infall signatures and a CO depletion factor) are not known, we will continue to refer to ‘Pers326’ as starless, not prestellar. The coordinates of Pers326 also coincides with an ammonia (NH3) peak (Rosolowsky et al., 2008; Kerr et al., 2019) and thus provides additional information such as the core’s kinetic temperature ($T_{\mathrm{k}}\sim 12$ K).
This criterion outlined above (Herschel plus NH3 observations) led to the investigation of the complex chemical inventory of Pers326. It was included in a sample of 35 total starless and prestellar cores in the Perseus Molecular Cloud surveyed with the Arizona Radio Observatory (ARO) 12 m dish, presenting high abundances of the simple complex organic molecules (COMs) CH3OH and CH3CHO (Scibelli et al., 2024). Follow-up observations were done with the Yebes Observatory 40 m telescope’s Q-band receiver to search for higher complexity COMs in a sub-sample of the 15 cores with bright CH3OH and CH3CHO, including Pers326 (also presented in Scibelli et al. 2024). The wide-band nature of the Q-band observations led to the serendipitous detection of the PN(1-0) line, concurrent with bright SiO(1-0) emission, which we present here and confirm with additional observations at higher frequencies.
@@ 2 Observations
For both the Yebes 40 m and IRAM 30 m datasets the reduction was done with the publicly available Python-based scripts [^1] https://github.com/andresmegias/gildas-class-python/ developed by Megías et al. (2023), which invokes the CLASS program of the GILDAS package (Pety 2005; Gildas Team 2013) for the combining of the spectra (again, see Scibelli et al. 2024 for more detail). Data were inspected, baselined, combined, and put on the main beam temperature, $T_{\mathrm{mb}}$, scale based on the public efficiency measurements given for each respective telescope and receiver. Gaussian fitting of spectral lines was done with the Python ‘Pyspeckit’ package (Ginsburg & Mirocha, 2011; Ginsburg et al., 2022). The lines presented and analyzed here are listed in Table 1 and plotted in Figures 2, 3, 4 and 5.
@@ 3 Detections and Line Properties
The wide-bandwidth in the Yebes 40 m data led to the serendipitous detection of the PN(1-0) line at 46.990 GHz toward the Pers326 starless core. A total of 15 cores bright in CH3OH and CH3CHO were observed in this sample (Scibelli et al., 2024), and so PN was also searched for in these cores with no success (see Table 2). Pers326 is the only core with strong ($T_{\mathrm{mb}}$ $>$ 1 K) SiO(1-0) emission at 43.423855 GHz [^2] Note that if one were to use the measured frequency of 43.42376 GHz from the JPL line list within the SPLATALOGUE database, SiO would be blue-shifted from the PN and PO emission by $\sim$0.8 km/s. This JPL entry has a roughly order of magnitude larger error than the entry in SLAIM, i.e., 50 kHz versus 3 kHz, respectively. One should thus be careful when extracting rest frequencies from SPLATALOGUE, and use those entries with smaller associated errors.  (Figure 2). This emission is only slightly blue-shifted, to 7.6 km/s, from the PN(1-0) $v_{\mathrm{lsr}}$ of 7.8 km/s, but given the coarse resolution of Yebes at these frequencies ($\sim$0.2 km/s) we conclude they are consistent (Table 1). A few other starless and prestellar cores reside nearby spatially to Pers326 with SiO detections, Pers264, Pers317 and Pers321 (see Figure 1), but the emission is much weaker ($T_{\mathrm{mb}}$ $<$ 0.3 K) and offset in $v_{\mathrm{lsr}}$ (Figure 2). For Pers264 the SiO(1-0) is blue-shifted more dramatically, to $\sim$2 km/s, suggesting this emission is not associated with the core itself (at a similar $v_{\mathrm{lsr}}$ to Pers326) but with the larger-scale surrounding SiO gas (see Figure 2). Given the strong SiO(1-0) and PN(1-0) correlation in Pers326, we suggest the origin of this gas-phase PN is related to shock chemistry (see section 5 for more discussion).
@@
Figure 2: SiO(1-0) emission observed toward Pers326 with the Yebes 40 m telescope (black) and a Gaussian fit to the line (blue). The vertical dashed lines correspond to the $v_{\mathrm{lsr}}$ of 7.6 km/s. Additional SiO emission toward the other nearby starless cores Pers321 (orange), Pers317 (red) and Pers264 (gray) are also plotted for reference.
@@
Figure 3: PN emission centered on the $F$ = 2-1 hyperfine component for each of the $J$ transitions observed toward Pers326 with both the Yebes 40 m and IRAM 30 m. As solid and blue and red curves are the Gaussian fits to the lines, the vertical dashed line shows a $v_{\mathrm{lsr}}$ of 7.6 km/s, and the horizontal gray line shows $1\sigma$ noise level. Spectra are offset by 100 mK for easier viewing. Annotated on the spectra are also the locations of the weaker and blended hyperfine components.
@@
Figure 4: PO emission observed toward Pers326 with the IRAM 30 m using both the 3 mm and 2 mm receivers. In solid red are the Gaussian fits to the detected ($>3\sigma$) lines and in dotted red are LTE fits at $T_{\mathrm{ex}}$ = 4.6 K. The vertical dashed line shows the consistent $v_{\mathrm{lsr}}$ of 7.6 km/s, and the horizontal gray line shows $1\sigma$ noise level. Spectra are offset by 80 mK for easier viewing. Full quantum numbers for each transition are given in Table 1.
@@
Figure 5: PO+ transitions observed toward Pers326. In solid red is the Gaussian fit to the detected ($>3\sigma$) line and the dotted lines are LTE fits at $T_{\mathrm{ex}}$ = 4.6 K. The vertical dashed line shows the consistent $v_{\mathrm{lsr}}$ of 7.6 km/s, and the horizontal gray line shows $1\sigma$ noise level. Spectra are offset by 40 mK for easier viewing.
@@ 4 Column Densities
Figure 6 shows, firstly, a least-squares fit to the PN data where $T_{\mathrm{ex}}=7.6\pm$0.7 K and $N_{\mathrm{tot}}=0.84\pm 0.12\times 10^{12}$ cm-2. A separate fit to the PO data finds a steeper slope where $T_{\mathrm{ex}}=4.63\pm$0.13 K and $N_{\mathrm{tot}}=2.63\pm 0.16\times 10^{12}$ cm-2. Uncertainties in the fits were calculated based on Monte Carlo simulations sampling the data points as Gaussians and then analyzing the resulting distributions of the parameters using the richvalues Python package [^3] https://github.com/andresmegias/richvalues. While the spread in $E_{\mathrm{u}}$ for PO is limited, with only two distinct values (8.4 K and 15.7 K), we also fit synthetic spectra to each individual hyperfine line to find the parameters derived from the rotation diagram fit agree well with the data (see Figure 4). Similarly, using the calculated $T_{\mathrm{ex}}$ value of 4.6 K for PO, we fit synthetic spectra to the PO+ transitions and find a column density roughly two orders of magnitude lower than for PO, at $N_{\mathrm{tot}}=3.02^{+0.15}_{-0.12}\times 10^{10}$ cm-2. All column densities are reported in Table 3.
As seen before in the literature (e.g., Fontani et al. 2016; Wurmser & Bergner 2022) this emission is expected to be subthermally excited. Future work to model these P-bearing species in the optically thick limit and with non-LTE codes in conjunction with a better understanding of the physical structure (e.g., volume density) of the emitting gas can shed light on this, but is beyond the scope for this current Letter.
Additionally, we note that it is still unclear if the PN, PO and PO+ emission comes from a compact region around the core or if they trace the larger-scale SiO emission. Therefore, we assume the emission fills our beams ($f=1$) in all cases and thus our $N$ values are likely lower limits. We do note that for PN if we assume a smaller filling fraction (using the same source size of 51.5 arcsec as found for methanol in Scibelli et al. 2024), the $T_{\mathrm{ex}}$ drops to 6.1 K and the total column density is slightly raised to $N_{\mathrm{tot}}=1.2\times 10^{12}$ cm-2.
@@ 5 Results and Discussion
We report the first detection of P-bearing molecules toward a shocked low-mass starless core. These observations provide new insight into the interstellar phosphorus budget in the earliest stage of star formation.
@@ 6 Conclusions
We thank the anonymous reviewer for their informative comments and feedback. The authors also thank Anthony Remijan and Víctor Rivilla for helpful discussion and suggestions. S.S. acknowledges the National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. I.J-.S and A.M. acknowledge funding from grant PID2022-136814NB-I00 funded by the Spanish Ministry of Science, Innovation and Universities/State Agency of Research MICIU/AEI/ 10.13039/501100011033 and by “ERDF/EU”, and from the ERC grant OPENS (project number 101125858) funded by the European Union. Y.S. was supported by National Science foundation Astronomy and Astrophysics grant AST-2205474. R.T.G. thanks the National Science Foundation for funding through the Astronomy & Astrophysics program (grant number 2206516). Observations carried out with the Yebes 40 m telescope (22A022 and 23A025). Paula Tarrío and Alba Vidal García carried out the observations and the first inspection of the data quality. The 40 m radio telescope at Yebes Observatory is operated by the Spanish Geographic Institute (IGN; Ministerio de Transportes, Movilidad y Agenda Urbana). Yebes Observatory thanks the European Union’s Horizon 2020 research and innovation programme for funding support to ORP project under grant agreement No 101004719. This work is also based on observations carried out under project number 024-24 with the IRAM 30 m telescope. IRAM is supported by INSU/CNRS (France), MPG (Germany) and IGN (Spain).